745 views

1 Answer

Best answer
2 2 votes
Here $\text{mean} (\beta) = 100$

so $\lambda = \large \frac{1}{\beta} =  \large \frac{1}{100}$

$P(X \leq x) = 1 - \large  e^{- \lambda x}$

$P(X \leq 200) = 1 - e^{-2} $

$P(X \geq 200) = 1 - P(X \leq 200)  $

$= 1 - (1 - e^{-2}) = e^{-2}  \approx   0.1353 $

Probability the bulb lasts more than $200$ hours is $0.1353$

$P(X \leq 50) = 1 - e^{- \frac{1}{100} \times 50}  \approx 0.3935$

Probability the bulb lasts less than $50$ hours is $0.3935$
• edited by
Position:
Show:

Related questions

0 0 votes
0 0 answers
607
607 views
Pooja Khatri asked Sep 27, 2018
607 views
Roll a die $n$ times and let $M$ be the number of times you roll $6$. Assume that $n$ is large.(c) How large should $n$ be so that the probability in(b) is larger than $0...
1 1 vote
1 1 answer
1.1k
1.1k views
Pooja Khatri asked Sep 27, 2018
1,102 views
Roll a die $n$ times and let $M$ be the number of times you roll $6$. Assume that $n$ is large.(b) Write down an approximation, in terms on $n$ and $\phi$, of the probabi...
0 0 votes
1 answers 1 answer
1.0k
1.0k views
Pooja Khatri asked Sep 27, 2018
1,001 views
Roll a die $n$ times and let $M$ be the number of times you roll $6$. Assume that $n$ is large.(a) Compute the expectation $EM$.
1 1 vote
0 0 answers
480
480 views
Pooja Khatri asked Sep 27, 2018
480 views
Toss a fair coin twice. You win $1$ dollar if at least one of the two tosses comes out heads.(b) Approximately how many times do you need to play so that you win at least...